# JEL Classifications: C61; D15; D52; E21; G51

### v2

@@ Appendix B Proof of Section 2 Results
$\hat{\beta}\hat{R}u^{\prime}[c(\hat{R}\left(w-\xi\right)+\hat{Y},\hat{Z}),\hat{Z}]\leq\hat{\beta}\hat{R}u^{\prime}[c(\hat{Y},\hat{Z}),\hat{Z}]\leq\hat{\beta}\hat{R}u^{\prime}(\hat{Y},\hat{Z})+\hat{\beta}\hat{R}M$ (24)
@@
[…]
@@
$\kappa=\max\left\{\mathbbm{E}_{z}\,\hat{\beta}\hat{R}\,h(\hat{R}\,[w-(u^{\prime}(\cdot,z))^{-1}(\kappa)]+\hat{Y},\hat{Z}),u^{\prime}(w,z)\right\},$ (25)
@@
$u^{\prime}(\xi,z)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\left(w-\xi\right)+\hat{Y},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\right\}.$ (26)
@@
To prove Theorem 2.1, we apply the Perov contraction theorem following Toda (2021b). For $h,g\in\mathscr{H}$ and $z\in\mathsf{Z}$, define
@@
Note that $d_{z}$ is always finite by the definition of $\mathscr{H}$. Define the vector-valued metric $d:\mathscr{H}\times\mathscr{H}\to\mathbbm{R}_{+}^{\mathsf{Z}}$ by $d(h,g)=(d_{1}(h,g),\dots,d_{Z}(h,g))$. By Proposition B.1 and the discussion in Section 2 of Toda (2021b), $(\mathscr{H},d)$ is a complete vector-valued metric space.
@@
Let $B:=K(1)$. Then $\tilde{T}$$\mathscr{H}\to\mathscr{H}$ is a Perov contraction with coefficient matrix $B$, that is, for any $h_{1},h_{2}\in\mathscr{H}$, we have
$d(\tilde{T}h_{1},\tilde{T}h_{2})\leq Bd(h_{1},h_{2}).$
@@
By Assumption 2.2 (2), we have $r(B)\in[0,1)$. By Theorem 3 of Toda (2021b), which is a generalization of the sufficient condition for contraction of Blackwell (1965), it suffices to verify the following two conditions:
@@
By Lemma B.2 and Theorem 2 of Toda (2021b), there exists a unique fixed point $h^{*}\in\mathscr{H}$ of $\tilde{T}$. From the topological conjugacy result in Lemma B.1, we have $\tilde{T}=HTH^{-1}$, so there exists a unique fixed point $c^{*}\in\mathscr{C}$ of $T$, and claim (i) is verified.
@@
$u^{\prime}(\xi(w,z),z)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\,[w-\xi(w,z)]+\hat{Y},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\right\}.$ (27)
@@
$\bar{w}_{c}(z):=u^{\prime}(\cdot,z)^{-1}\left[\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{Y},\hat{Z}),\hat{Z})\right]\quad\text{and}\quad\bar{w}(z):=\bar{w}_{c^{*}}(z).$ (28)
@@
[…]
$\leq\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{Y},\hat{Z}),\hat{Z})=u^{\prime}(\bar{w}_{c}(z),z).$
@@
To establish our key results, we extend the operator $F_{i}$ defined in (15) as follows. With slight abuse of notation, in what follows, for each $i\in\{1,\dots,N\}$ and $j\in\{i,\dots,M\}$, we define $F_{i}:[1,\infty]^{M}\to[1,\infty]^{M}$ by
$(F_{i}x)(\tilde{z}_{j}):=\begin{cases}\infty,&\text{if\, $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$},\\ \left(1+(G_{i}x)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}},&\text{if\, $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$}.\end{cases}$ (29)
@@
$\limsup_{w\to\infty}\frac{Tc(w,z_{ij})}{w}\leq(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$ (30)
@@
$\limsup_{n\to\infty}\lambda_{n}(z_{ij})\leq x(\hat{Z})^{-1/\gamma(\hat{Z})}\hat{R}(1-\alpha(z_{ij})).$ (32)
@@
$\lambda_{n}(z_{ij})\leq\hat{R}(1-\alpha_{n}(z_{ij}))+\frac{\hat{Y}}{w_{n}}\to\hat{R}(1-\alpha(z_{ij}))=0,$
@@
$\alpha_{n}(z_{ij})^{-\gamma_{i}}\geq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$. (33)
Case 1. If $\sum_{j=1}^{i-1}\bar{p}_{ij}>0$, then there exists at least one $k<i$ such that $\bar{p}_{ik}>0$. By definition, this implies that $\gamma_{i}-\gamma_{k}>0$ with positive probability. Letting $n\to\infty$ in (33) and applying Fatou’s lemma, we have
@@
$\alpha(z_{ij})^{-\gamma_{i}}$ $\geq\bar{p}_{ik}\tilde{p}_{j\ell}\mathbbm{E}\left(\hat{\beta}\hat{R}\big[x(z_{k\ell})^{-1/\gamma_{k}}\hat{R}(1-\alpha(z_{ij}))\big]^{-\gamma_{k}}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma_{k}}\right)=\infty,$
where the expectation $\mathbbm{E}$ is taken with respect to the innovation $\hat{\varepsilon}$, and the last equality holds because, in this case, $\mathbbm{P}_{(z,\hat{z})}(\hat{\beta}\hat{R}>0)>0$ for all $(z,\hat{z})\in\mathsf{Z}^{2}$ by assumption. In addition, since by definition $(F_{i}x)(\tilde{z}_{j})\equiv\infty$ in this case, we have
@@
Case 2. If $\sum_{j=1}^{i-1}\bar{p}_{ij}=0$ and $\bar{p}_{ii}>0$, then, considering only the transition of state $\bar{z}_{i}$ to itself next period, (33) implies that
@@
For given $i\in\{1,\dots,N\}$, if $\sum_{j=1}^{i-1}\bar{p}_{ij}=0$ and $\bar{p}_{ii}>0$, then $F_{i}$ defined in (29) has a fixed point $x_{i}^{*}\in\mathbbm{R}_{+}^{M}$ if and only if $r(G_{i})<1$, and the fixed point is unique. Take any $x_{0}\in\mathbbm{R}_{+}^{M}$, and define the sequence $\{x_{n}\}_{n=1}^{\infty}\subset\mathbbm{R}_{+}^{M}$ by $x_{n}=F_{i}x_{n-1}$ for all $n\in\mathbb{N}$. Then we have the following:
@@
[…]
@@
$0\leq\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq\limsup_{w\to\infty}\frac{c_{n}(w,z_{ij})}{w}\leq x_{n}(\tilde{z}_{j})^{-1/\gamma_{i}}.$ (34)
@@
$\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$ (35)
@@
$\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\geq x^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}$ (36)
@@
$\lim_{w\to\infty}\frac{Tc(w,z_{ij})}{w}=(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$ (37)
@@
$\lim_{n\to\infty}\lambda_{n}(z_{ij})=x(\hat{Z})^{-1/\gamma(\hat{Z})}\hat{R}(1-\alpha(z_{ij})).$ (38)
@@
$=\max\left\{\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\bigg[\frac{c(\hat{R}(w_{n}-\alpha_{n}(z_{ij})w_{n})+\hat{Y},\hat{Z})}{w_{n}}\bigg]^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})},\;1\right\}.$ (39)
@@
$\liminf_{n\to\infty}\lambda_{n}(z_{ij})\geq x(\hat{Z})^{-1/\gamma(\hat{Z})}m(1-\alpha(z_{ij})).$ (40)
@@
$\alpha_{n}(z_{ij})^{-\gamma_{i}}=\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\bigg[\frac{c(\hat{R}(w_{n}-\alpha_{n}(z_{ij})w_{n})+\hat{Y},\hat{Z})}{w_{n}}\bigg]^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$. (41)
@@
there exists $N\in\mathbb{N}$ such that $\lambda_{n}\geq\underline{\lambda}$, $\alpha_{n}<\bar{\alpha}$, and $w_{n}\geq 1$ for all $n\geq N$ and $z_{ij}$. The integrand in the expectation in (39) is bounded above by
$\max_{z,\hat{z}\in\mathsf{Z}}\hat{\beta}\hat{R}\underline{\lambda}^{-\gamma(\hat{z})}$$\infty$,
which is integrable. Letting $n\to\infty$, applying the dominated convergence theorem, and using $\alpha(z_{ij})<1$, we obtain
$\alpha(z_{ij})^{-\gamma_{i}}$ $=\lim_{n\to\infty}\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$
$=\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma(\hat{Z})}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$.
@@
$\alpha(z_{ij})^{-\gamma_{i}}$ $=\bar{p}_{ii}\sum_{h=1}^{M}\tilde{p}_{jh}\mathbbm{E}_{z_{ij},z_{ih}}\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma_{i}}$
$+\sum_{k=i+1}^{N}\bar{p}_{ik}\sum_{h=1}^{M}\tilde{p}_{jh}\mathbbm{E}_{z_{ij},z_{ih}}\left(\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma_{k}}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma_{k}}\right)$
$=\bar{p}_{ii}\sum_{h=1}^{M}\tilde{p}_{jh}\mathbbm{E}_{z_{ij},z_{ih}}\hat{\beta}\hat{R}\left(x(\tilde{z}_{h})^{-1/\gamma_{i}}\hat{R}(1-\alpha(z_{ij}))\right)^{-\gamma_{i}},$
@@
$\lim_{w\to\infty}\frac{Tc(w,z_{ij})}{w}$ $=\alpha(z_{ij})=\frac{1}{1+\left(\bar{p}_{ii}\sum_{h=1}^{M}\tilde{p}_{jh}\mathbbm{E}_{z_{ij},z_{ih}}\hat{\beta}\hat{R}^{1-\gamma_{i}}x(\tilde{z}_{h})\right)^{1/\gamma_{i}}}$
@@
$c^{*}(w,z)<w(1-\varepsilon)\quad\Longrightarrow\quad\frac{c^{*}(w,z)}{w}<1-\varepsilon.$
@@
$\frac{c^{*}(w^{\prime},z)}{w^{\prime}}\geq 1-\varepsilon.$
@@
$\frac{c^{*}(w,z)}{w}<\frac{c^{*}(w^{\prime},z)}{w^{\prime}}$
@@
$(J_{i}y)(\tilde{z}_{j}):=\left(1+(U_{i}y)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}}.$ (42)
@@
$\lim_{w\to\infty}\frac{c^{*}(w,z_{ij};\psi,\gamma_{i})}{w}=y_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$ (43)
@@
Fix $i\in\{1,\dots,N\}$. Suppose $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and there exists $b>0$ such that $Y(z,\hat{z},\hat{\varepsilon})\geq b$ almost surely conditional on (z,z^)($\hat{z}$). Choose $\psi$ such that
$b^{-\gamma_{h}}/\psi_{h}\leq b^{-\gamma_{i}}/\psi_{i}$
for all $h\geq i$. Then $c^{*}(w,z_{ij};\psi,\gamma_{i})\leq c^{*}(w,z_{ij})$ for all $w\geq b$ and $z_{ij}\in\mathsf{Z}$.
@@
Denote $c_{n}^{0}:=T_{0}^{n}c$ and $c_{n}:=T^{n}c$. The claim obviously holds for $n=0$. Suppose it holds for arbitrary $n$. Then $c_{n}^{0}(w,z_{ij})\leq c_{n}(w,z_{ij})$. It remains to verify that $c_{n+1}^{0}(w,z_{ij})\leq c_{n+1}(w,z_{ij})$ for all $w\geq b$ and $z_{ij}\in\mathsf{Z}$. Suppose this is not true. Then there exists $w\geq b$ and $z_{ij}\in\mathsf{Z}$ such that $c_{n+1}^{0}(w,z_{ij})>c_{n+1}(w,z_{ij})$. Then we have
$1\geq\frac{c_{n+1}^{0}(w,z_{ij})}{w}>\frac{c_{n+1}(w,z_{ij})}{w}.$
By the definition of $T$ and the induction argument, we have
@@
$\leq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}c_{n}^{0}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},\hat{Z})^{-\hat{\gamma}}$. (44)

### v3

@@ Appendix B Proof of Section 2 Results
(23) $\hat{\beta}\hat{R}u^{\prime}[c(\hat{R}\left(w-\xi\right)+\hat{Y},\hat{Z}),\hat{Z}]\leq\hat{\beta}\hat{R}u^{\prime}[c(\hat{Y},\hat{Z}),\hat{Z}]\leq\hat{\beta}\hat{R}u^{\prime}(\hat{Y},\hat{Z})+\hat{\beta}\hat{R}M.$
@@
and the correspondence $(w,z)\mapsto I(w,z)$ is nonempty, compact-valued and continuous, Theorem B.1.4 of Stachurski (2009) implies that $Tc$ is continuous on $S\SS$.
@@
(24) $\kappa=\max\left\{\mathbbm{E}_{z}\,\hat{\beta}\hat{R}\,h(\hat{R}\,[w-(u^{\prime}(\cdot,z))^{-1}(\kappa)]+\hat{Y},\hat{Z}),u^{\prime}(w,z)\right\},$
@@
(25) $u^{\prime}(\xi,z)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\left(w-\xi\right)+\hat{Y},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\right\}.$
@@
To prove Theorem 2.1, we apply the Perov contraction theorem following Toda (2021). For $h,g\in\mathscr{H}$ and $z\in\mathsf{Z}$, define
@@
Note that $d_{z}$ is always finite by the definition of $\mathscr{H}$. Define the vector-valued metric $d:\mathscr{H}\times\mathscr{H}\to\mathbbm{R}_{+}^{\mathsf{Z}}$ by $d(h,g)=(d_{1}(h,g),\dots,d_{Z}(h,g))$. By Proposition B.1 and the discussion in Section 2 of Toda (2021), $(\mathscr{H},d)$ is a complete vector-valued metric space.
@@
Let $B:=K(1)$. Then $\tilde{T}$$\mathscr{H}\to\mathscr{H}$ is a Perov contraction with coefficient matrix $B$, that is, for any $h_{1},h_{2}\in\mathscr{H}$, we have $d(\tilde{T}h_{1},\tilde{T}h_{2})\leq Bd(h_{1},h_{2})$.
@@
By Assumption 2.2 (2), we have $r(B)\in[0,1)$. By Theorem 3 of Toda (2021), it suffices to verify the following two conditions:
@@
By Lemma B.2 and Theorem 2 of Toda (2021), there exists a unique fixed point $h^{*}\in\mathscr{H}$ of $\tilde{T}$. From the topological conjugacy result in Lemma B.1, we have $\tilde{T}=HTH^{-1}$, so there exists a unique fixed point $c^{*}\in\mathscr{C}$ of $T$, and claim (i) is verified.
@@
(26) $u^{\prime}(\xi(w,z),z)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\,[w-\xi(w,z)]+\hat{Y},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\right\}.$
@@
(27) $\bar{w}_{c}(z):=u^{\prime}(\cdot,z)^{-1}\left[\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{Y},\hat{Z}),\hat{Z})\right]\quad\text{and}\quad\bar{w}(z):=\bar{w}_{c^{*}}(z).$
@@
$u^{\prime}(w,z)<\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\left[w-\xi(w,z)\right]+\hat{Y},\hat{Z}),\hat{Z})\leq\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{Y},\hat{Z}),\hat{Z})=u^{\prime}(\bar{w}_{c}(z),z).$
@@
To establish our key results, we extend the operator $F_{i}$ defined in (15) as follows. With slight abuse of notation, in what follows, for each $i\in\{1,\dots,N\}$ and $j\in\{1,\dots,M\}$, we define $F_{i}:[1,\infty]^{M}\to[1,\infty]^{M}$ by
(28) $(F_{i}x)(\tilde{z}_{j}):=\begin{cases}\infty,&\text{if\, $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$},\\ \left(1+(G_{i}x)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}},&\text{if\, $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$}.\end{cases}$
@@
(29) $\limsup_{w\to\infty}\frac{Tc(w,z_{ij})}{w}\leq(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$
@@
(31) $\limsup_{n\to\infty}\lambda_{n}(z_{ij})\leq x(\hat{Z})^{-1/\gamma(\hat{Z})}\hat{R}(1-\alpha(z_{ij})).$
@@
$\lambda_{n}(z_{ij})\leq\hat{R}(1-\alpha_{n}(z_{ij}))+\hat{Y}/w_{n}\to\hat{R}(1-\alpha(z_{ij}))=0,$
@@
(32) $\alpha_{n}(z_{ij})^{-\gamma_{i}}\geq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$.
Case 1. If $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$, then there exists at least one $h<i$ such that $\bar{p}_{ih}>0$. By definition, this implies that $\gamma_{i}-\gamma_{h}>0$ with positive probability. Letting $n\to\infty$ in (32) and applying Fatou’s lemma, we have
@@
$\alpha(z_{ij})^{-\gamma_{i}}$ $\geq\bar{p}_{ih}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{hk}}\left(\hat{\beta}\hat{R}\big[x(z_{hk})^{-1/\gamma_{h}}\hat{R}(1-\alpha(z_{ij}))\big]^{-\gamma_{h}}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma_{h}}\right)=\infty,$
where the expectation $\mathbbm{E}_{z_{ij},z_{hk}}$ is taken with respect to the innovation $\hat{\varepsilon}$, and the last equality holds because, in this case, $\mathbbm{P}_{(z,\hat{z})}(\hat{\beta}\hat{R}>0)>0$ for all $(z,\hat{z})\in\mathsf{Z}^{2}$ by assumption. In addition, since by definition $(F_{i}x)(\tilde{z}_{j})\equiv\infty$ in this case, we have
@@
Case 2. If $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and $\bar{p}_{ii}>0$, then, considering only the transition of state $\bar{z}_{i}$ to itself next period, (32) implies that
@@
For given $i\in\{1,\dots,N\}$, if $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and $\bar{p}_{ii}>0$, then $F_{i}$ defined in (28) has a fixed point $x_{i}^{*}\in\mathbbm{R}_{+}^{M}$ if and only if $r(G_{i})<1$, and the fixed point is unique. Take any $x_{0}\in\mathbbm{R}_{+}^{M}$, and define the sequence $\{x_{n}\}_{n=1}^{\infty}\subset\mathbbm{R}_{+}^{M}$ by $x_{n}=F_{i}x_{n-1}$ for all $n\in\mathbb{N}$. Then we have the following:
@@
Define the sequence $\{c_{n}\}\subset\mathscr{C}$ by $c_{0}(w,z)=w$ and $c_{n}(w,z)=Tc_{n-1}(w,z)$ for all $n\geq 1$. Since $Tc(w,z)\leq w$ for any $c\in\mathscr{C}$, we have $c_{1}(w,z)=Tc_{0}(w,z)\leq w=c_{0}(w,z)$. Since $T:\mathscr{C}\to\mathscr{C}$ is monotone, by induction $0\leq c_{n}(w,z)\leq c_{n-1}(w,z)\leq w$ for all $n\geq 1$ $S(w,z)\in\SS, and$ $c^{*}(w,z)=\lim_{n\to\infty}c_{n}(w,z)$ exists. This $c^{*}$ is the unique fixed point of $T$ and also the unique optimal policy in $\mathscr{C}$.
@@
(33) $0\leq\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq\limsup_{w\to\infty}\frac{c_{n}(w,z_{ij})}{w}\leq x_{n}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
@@
(34) $\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
@@
(35) $\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\geq x^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}$
@@
(36) $\lim_{w\to\infty}\frac{Tc(w,z_{ij})}{w}=(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$
@@
(37) $\lim_{n\to\infty}\lambda_{n}(z_{ij})=x(\hat{Z})^{-1/\gamma(\hat{Z})}\hat{R}(1-\alpha(z_{ij})).$
@@
(38) $=\max\left\{\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\bigg[\frac{c(\hat{R}(w_{n}-\alpha_{n}(z_{ij})w_{n})+\hat{Y},\hat{Z})}{w_{n}}\bigg]^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})},\;1\right\}.$
@@
(39) $\liminf_{n\to\infty}\lambda_{n}(z_{ij})\geq x(\hat{Z})^{-1/\gamma(\hat{Z})}m(1-\alpha(z_{ij})).$
@@
(40) $\alpha_{n}(z_{ij})^{-\gamma_{i}}=\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\bigg[\frac{c(\hat{R}(w_{n}-\alpha_{n}(z_{ij})w_{n})+\hat{Y},\hat{Z})}{w_{n}}\bigg]^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$.
@@
there exists $N\in\mathbb{N}$ such that $\lambda_{n}\geq\underline{\lambda}$, $\alpha_{n}<\bar{\alpha}$, and $w_{n}\geq 1$ for all $n\geq N$ and $z_{ij}$. The integrand in the expectation in (38) is bounded above by $\max_{z,\hat{z}\in\mathsf{Z}}\hat{\beta}\hat{R}\underline{\lambda}^{-\gamma(\hat{z})}$$\infty$, which is integrable. Letting $n\to\infty$, applying the dominated convergence theorem, and using $\alpha(z_{ij})<1$, we obtain
$\alpha(z_{ij})^{-\gamma_{i}}=\lim_{n\to\infty}\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}=\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma(\hat{Z})}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$.
@@
$\alpha(z_{ij})^{-\gamma_{i}}$ $=\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma_{i}}$
$+\sum_{h=i+1}^{N}\bar{p}_{ih}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{hk}}\left(\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma_{h}}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma_{h}}\right)$
$=\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}\left(x(\tilde{z}_{k})^{-1/\gamma_{i}}\hat{R}(1-\alpha(z_{ij}))\right)^{-\gamma_{i}},$
@@
$\lim_{w\to\infty}\frac{Tc(w,z_{ij})}{w}$ $=\alpha(z_{ij})=\frac{1}{1+\left(\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}^{1-\gamma_{i}}x(\tilde{z}_{k})\right)^{1/\gamma_{i}}}$
@@
$c^{*}(w,z)<w(1-\varepsilon)\quad\Longrightarrow\quad c^{*}(w,z)/w<1-\varepsilon.$
@@
$c^{*}(w^{\prime},z)/w^{\prime}\geq 1-\varepsilon.$
@@
$c^{*}(w,z)/w<c^{*}(w^{\prime},z)/w^{\prime}$
@@
(41) $(J_{i}y)(\tilde{z}_{j}):=\left(1+(U_{i}y)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}}.$
@@
(42) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij};\psi,\gamma_{i})}{w}=y_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
@@
Fix $i\in\{1,\dots,N\}$. Suppose $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and there exists $b>0$ such that $Y(z,\hat{z},\hat{\varepsilon})\geq b$ almost surely conditional on (z,z^)($\hat{z}$). Choose $\psi$ such that $b^{-\gamma_{h}}/\psi_{h}\leq b^{-\gamma_{i}}/\psi_{i}$ for all $h\geq i$. Then $c^{*}(w,z_{ij};\psi,\gamma_{i})\leq c^{*}(w,z_{ij})$ for all $w\geq b$ and $z_{ij}\in\mathsf{Z}$.
@@
Denote $c_{n}^{0}:=T_{0}^{n}c$ and $c_{n}:=T^{n}c$. The claim obviously holds for $n=0$. Suppose it holds for arbitrary $n$. Then $c_{n}^{0}(w,z_{ij})\leq c_{n}(w,z_{ij})$. It remains to verify that $c_{n+1}^{0}(w,z_{ij})\leq c_{n+1}(w,z_{ij})$ for all $w\geq b$ and $z_{ij}\in\mathsf{Z}$. Suppose this is not true. Then there exists $w\geq b$ and $z_{ij}\in\mathsf{Z}$ such that $c_{n+1}^{0}(w,z_{ij})>c_{n+1}(w,z_{ij})$. Then we have $1\geq c_{n+1}^{0}(w,z_{ij})/w>c_{n+1}(w,z_{ij})/w.$ By the definition of $T$ and the induction argument, we have
@@
(43) $\leq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}c_{n}^{0}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},\hat{Z})^{-\hat{\gamma}}$.